OpenStudy (anonymous):

89x=7(13x-8) my answer is x=28

OpenStudy (angela210793):

89x=91x-56 89x-91x=-56 -2x=-56 x=28 :D

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thanks